Solve by the elimination method.
9x + 3y = -39
4x - 3y = -26
Will someone help me with this one please?

Answer by oscargut(2103)

9x + 3y = -39
4x - 3y = -26
adding equations we have:
13x=-65
x=-65/13=-5
4(-5)-3y=-26
-3y=-6
y=2
Answer: x=-5,y=2